# Polarization fringes in optical systems: a compendium

### v1

@@ 3.3 Other considerations about fringe mitigation
    The suppression of fringes by averaging over the angular span of a converging or diverging beam will in general be accompanied by some degree of depolarization (see Sect. 5 and App. B of [3], and Ch. 6 of [22]). An example of this can be seen in Fig. 10, particularly in the $QQ/II$ element, which shows a clear reduction of the spectral mean value of that Mueller element going from a collimated beam (black curve) to an f/13 beam (orange curve). Depolarization can be modeled, measured, and accounted for in calibration algorithms if necessary [23, 25, 24, 26, 3].
@@ 4 Conclusions
While the modeling framework developed in this work in general does not allow us to reproduce the results of the Berreman calculus in all the details [4, 5, 3], nonetheless it fulfills the purpose of predicting with sufficient accuracy both the pattern and the magnitude of spectral and spatial polarization fringes, in order to estimate their impact on the performance of polarization modulators and achromatic retarders, using realistic instrument design parameters (such as spectral resolution and f/#; see [3]). It also provides a convenient and sufficiently reliable tool for the validation of optical-design strategies of fringe mitigation for different polarimetric applications and instrument concepts. To demonstrate this, we modeled various types of birefringent optics, and showed in detail how fringes can be manipulated and effectively suppressed in the observed data by making informed design choices.
@@
To conclude, we note that such a formalism also allows us to model the formation of polarization fringes directly in the observed data. This is achieved by adopting the “fringed” Mueller matrix of the system, averaged over the angular range of the incoming optical beam (see, e.g., Fig. 11), in the data simulation pipeline that maps the incoming Stokes vector from a polarized source to the modulated signals measured by a detector. These modulated signals are then converted back to Stokes profiles using a reference demodulation matrix, which is usually determined via polarization calibration under environmental conditions of the system similar to those of the observations. The product of this demodulation matrix with the “fringed” modulation matrix generally does not correspond to an identity, thus producing residual non-diagonal elements that account for polarization errors such as cross-talk. The demodulation of the signals produced by a “fringed” modulation matrix will naturally produce a fringe pattern that is superimposed to the true polarization signals from the source. This type of modeling can be used to test modulator designs specifically aimed at mitigating the effects of polarization fringes on Stokes measurements, and also to create synthetic observations with fringes that can be used to test the applicability and effectiveness of post-processing de-fringing methods [17].
5 Acknowledgments and Disclosures
@@
This material is based upon work supported by the National Center for Atmospheric Research, which is a major facility sponsored by the National Science Foundation (NSF) under Cooperative Agreement No. 1852977. DMH acknowledges support by the NSF Daniel K. Inouye Solar Telescope (DKIST) project. For the birefringence data of SiO2 and MgF2 used for the models in this work, we relied on the measurements by S. Sueoka and D. Elmore for the DKIST project [27].
The authors declare no conflicts of interest.

### v2

@@ 3.3 Other considerations about fringe mitigation
    The suppression of fringes by averaging over the angular span of a converging or diverging beam will in general be accompanied by some degree of depolarization (see Sect. 5 of [3], and Ch. 6 of [22]). Therefore, depolarization must be modeled during system design, to ensure amplitudes remain below the allocated error budget. Common error budgets in solar telescopes adopt depolarization limits near 1% (e.g. DKIST [21], Hinode [23]). It is also important to remark that depolarization can easily be measured in the lab to validate models (c.f. [3], App. B), and if necessary it can be accounted for in calibration algorithms by including appropriate variables in the system model [24, 26, 25, 27, 3].
@@ 4 Conclusions
The modeling framework developed in this work may not allow us to reproduce all numerical details of more rigorous treatments such as Berreman calculus [4, 8, 5, 22], but nonetheless it still fulfills the purpose of effectively predicting both pattern and magnitude of the polarization fringes produced in the presence of plane-parallel optics inside an instrument. The accuracy of our approximate treatment is completely adequate to quantitatively assess the impact of such fringes on the polarimetric performance of the instrument, based on realistic instrument design parameters, such as spectral resolution and f/# [3], and how those design parameters need to be modified in order to reduce the impact of fringes. It therefore provides a convenient and sufficiently reliable tool for the validation of optical-design strategies of fringe mitigation for different polarimetric applications and instrument concepts. To demonstrate this, we modeled various types of birefringent optics, and showed in detail how fringes can be manipulated and effectively suppressed in the observed data by making informed design choices. In fact, this type of modeling has already been used to guide the optical design of major nationally funded observing facilities, such as the DKIST [13, 15, 16], in order to quantify and mitigate the impact of these polarimetric artifacts on the science deliverables of the project.
@@
As a final note, we point out how this type of modeling also allows us to simulate the manifestation of polarization fringes directly in the science data. While we have not provided a practical example of this, we briefly describe the concept here. In order to realistically simulate the polarimetric performance of an instrument, one must model its polarization modulation matrix [12], properly averaged over the angular range of the incoming optical beam of the instrument and at the proper spectral sampling resolution (see, e.g., Fig. 11). When such a modulation matrix is adopted for the data simulation pipeline of the instrument, which maps the incoming Stokes vector from a polarized target to the modulated signals measured by the detector, the simulated signals will be affected by the presence of spatial and spectral fringes. On the other hand, the data reduction for a polarimetric instrument involves the use of a reference demodulation matrix that decodes the modulated signals into the Stokes profiles of the target. Such a demodulation matrix must be determined via polarization calibration of the instrument under environmental conditions that ideally should be identical to those of the observations. Because such an ideal occurrence is never realized in practice, and fringes are highly sensitive to even small variations of the environmental conditions, the product of the reference demodulation matrix with the actual modulation matrix generally does not correspond to an identity, the residuals being responsible for the introuction of polarization errors such as cross-talk [28] and FOV-dependent fringes. In other words, the “blind” demodulation of signals produced by a “fringed” modulation matrix will naturally cause the appearance of fringe patterns that are superimposed on the true polarization signals from the source. This type of modeling can be used to test modulator designs specifically aimed at mitigating the effects of polarization fringes on Stokes measurements, and also to create synthetic observations with fringes that can be used to test the applicability and effectiveness of post-processing de-fringing methods [17] or the resilience of Stokes inversion codes to the presence of fringes in the data.
We expect that such in-depth design analyses will become increasingly relevant for the development of future instrumentation, and envision that the use of modeling tools such as the one presented in this work will become commonplace in the design phase of any polarimetric instrument.
@@
Acknowledgments and disclosures. This material is based upon work supported by the NSF National Center for Atmospheric Research, which is a major facility sponsored by the National Science Foundation under Cooperative Agreement No. 1852977. DMH acknowledges support by the NSF Daniel K. Inouye Solar Telescope (DKIST) project. The authors declare no conflicts of interest.
Data availability. Data underlying the results presented in this paper were exclusively modeled. They are not publicly available at this time, but they may be obtained from the authors upon reasonable request. The birefringence data of SiO2 and MgF2 used for the models in this work rely on the measurements by S. Sueoka and D. Elmore for the DKIST project [29].